# Local Normalization Distortion and the Thermodynamic Formalism of Decoding Strategies for Large Language Models

### v1

@@ 7 Conclusions
By expressing popular decoding strategies as equilibrium states, one can see that the quantity which they maximize contains a term relating to local normalization of probability mass which seems unrelated to any reasonable goal of a decoding strategy. In particular, this term pulls the resulting probability distribution away from the quality-diversity maximizing curve. We have shown experimentally that the effect of local normalization distortion on the probability of selecting a string is typically very large (Section 6.1), and that it has a strongly negative effect on the quality-diversity tradeoff (Section 6.2). This accounts for the under-performance of top-k sampling relative to nucleus sampling on this metric. All of these factors lead us to the conclusion that local normalization distortion has a profoundly negative effect on machine-generated text and that it should be carefully considered both when practitioners choose a decoding strategy and in the design of future methods.
@@ 9 Limitations
Our experiments are run on the open-source Llama 2 language model. While this is not uncommon for research in computational linguistics, the setting in which decoding strategies such as temperature sampling are most widely deployed is in closed-source models such as ChatGPT OpenAI (2022). Our theoretical results hold for all language models and we do not believe the conclusions of our experimental section would change with language model. It is however the case that the magnitude of local normalization distortion would decrease if the entropies of the next token probability distributions were typically lower. Thus, language models with higher certainty about their next token predictions would give rise to smaller numbers in Table 1, for example.
We have used log-likelihood as a proxy for the quality of machine-generated text and entropy as a proxy for its diversity. These metrics are clearly imperfect. In particular, human judgement of text quality may be a better metric for the quality of a text, although obtaining these human judgements often prohibitive due to cost Clark et al. (2021).
@@ Appendix C Global Temperature Normalization in Ergodic Theory and Fractal Geometry
We mentioned in section 3.1 that global normalization of measures is the standard method in statistical physics, ergodic theory and fractal geometry. A short explanation of this remark is that globally normalized temperature sampling corresponds to taking the Gibbs-equilibrium measure associated to potential $\log p$ at temperature $\frac{1}{\tau}$. Similarly, when using a truncation sampling algorithm with allowed set $\mathcal{A}$, globally normalized sampling corresponds to sampling from the Gibbs-equilibrium measure associated to potential $\log p$ on the sequence space defined by $\mathcal{A}$. For both of these comments, see Bowen (2008).
@@
[…]
One might ask what can be said about the set of texts for which the average log probability of tokens takes some different value $\alpha$. How are typical such texts distributed? How many are there? Such questions are answered through the multifractal analysis of ergodic averages, see for example Falconer (2007). Solutions involve globally (rather than locally) normalized temperature sampling.
@@ Appendix D Proofs
We recall Proposition 3.1, which stated that the limit as temperature tends to zero of locally normalized temperature sampling is greedy decoding, whereas the limit as temperature tends to zero of globally normalized temperature sampling is the distribution achieving globally maximal average log likelihood.
@@
$=$ $\dfrac{p({\bf w|y_{<m}})}{\prod_{i=0}^{T-m}Z_{k}(y_{0}\cdots y_{m}w_{m+1}\cdots w_{m+i-1})}$
@@
Appendix E Supplementary figures
@@
Figure 3: For each graph we generate pairs of texts according to top-k, nucleus or temperature sampling. Parameters for the decoding strategies are tuned so that typical renormalizing quantitites $Z_{k},$ $Z_{\pi}$ and $Z_{\tau}$ are of the same size. We then plot the log of the ratio of the local normalization distortion for the pair of generated texts, and then plot the cumulative distribution function. We see in each case that nucleus sampling produces the smallest local normalization distortion, followed by temperature sampling and top-k sampling.
@@ Appendix F A Comment on Timescales for Normalization
$\frac{1}{C}\leq\dfrac{q_{\tau}^{\prime\prime}(\{\underline{z}\in\mathcal{A}^{\infty}:z_{1}\cdots z_{T}=w_{1}\cdots w_{T}\})}{p^{\frac{1}{\tau}}(w_{1}\cdots w_{T})\exp(T.P)}\leq C.$
@@
Appendix G Licenses
We have used Llama 2-7B under the community license agreement, see https://huggingface.co/meta-llama/Llama-2-7b. Use for research is consistent with the terms of this license.

### v2

@@ 7 Conclusions
Our primary contribution is to express popular decoding strategies as equilibrium states. In doing so, one can see that the quantity which they maximize contains a term relating to local normalization of probability mass which seems unrelated to any reasonable goal of a decoding strategy. In particular, this term pulls the resulting probability distribution away from the quality-diversity maximizing curve. We have shown experimentally that the effect of local normalization distortion on the probability of selecting a string is typically very large (Section 6.1), and that it has a strongly negative effect on the quality-diversity tradeoff (Section 6.2) when these quantities are measured through entropy and log-likelihood. In a follow-up work Kempton et al. (2025) we build a detector of machine generated text based on local normalization distortion which outperforms state of the art alternatives. These factors lead us to the conclusion that local normalization distortion may have a negative effect on machine-generated text and that it should be carefully considered both when practitioners choose a decoding strategy and in the design of future methods for detecting machine-generated text.
@@ 9 Limitations
Our experiments are run on the open-source Llama 2 language model (Touvron et al., 2023). While this is not uncommon for research in computational linguistics, the setting in which decoding strategies such as temperature sampling are most widely deployed is closed-source models such as with ChatGPT OpenAI (2022). Our theoretical results hold for all language models and we do not believe the conclusions of our experimental section would change with language model. It is however the case that the magnitude of local normalization distortion would decrease if the entropies of the next token probability distributions were typically lower. Thus, language models with higher certainty about their next token predictions would give rise to smaller numbers in Table 1, for example.
We have used log-likelihood as a proxy for the quality of machine-generated text and entropy as a proxy for its diversity. These metrics are clearly imperfect. In particular, human judgement of text quality may be a better metric for the quality of a text, although obtaining these human judgments is often prohibitive due to cost Clark et al. (2021).
@@ Appendix B Rejection sampling algorithms
For further information on rejection sampling and faster algorithms which approximate rejection sampling see Lipkin et al. (2025).
@@ Appendix C Global Temperature Normalization in Ergodic Theory and Fractal Geometry
We mentioned in Section 3.1 that global normalization of measures is a standard method in statistical physics, ergodic theory and fractal geometry. A short explanation of this remark is that globally normalized temperature sampling corresponds to taking the Gibbs-equilibrium measure associated to potential $\log p$ at temperature $\frac{1}{\tau}$. Similarly, when using a truncation sampling algorithm with allowed set $\mathcal{A}$, globally normalized sampling corresponds to sampling from the Gibbs-equilibrium measure associated to potential $\log p$ on the sequence space defined by $\mathcal{A}$. For both of these comments, see Bowen (2008).
@@
$\int_{\mbox{contexts }v_{<t}}\int_{v\in\mathcal{V}}p(v|v_{<t})dp(v|v_{<t})dp(v_{<t}).$
One might ask what can be said about the set of texts for which the average log probability of tokens takes some different value $\alpha$. How are typical such texts distributed? How many are there? Such questions are answered through the multifractal analysis of ergodic averages, see, for example, Falconer (2007). Solutions involve globally (rather than locally) normalized temperature sampling.
@@ Appendix D Proofs
We recall Proposition 3.1, which stated that the limit as temperature tends to zero of locally normalized temperature sampling is greedy decoding, whereas the limit as temperature tends to zero of globally normalized temperature sampling is the distribution achieving globally maximal average log-likelihood.
@@
$=$ $\dfrac{p({\bf w|y_{<m}})}{\prod_{i=0}^{T-m}Z_{k}(y_{0}\cdots y_{m}w_{m+1}\cdots w_{m+i-1})}.$
@@
Appendix E Supplementary Figures and Replication on other Language Models
@@
Figure 3: For each graph we generate pairs of texts according to top-k, nucleus or temperature sampling with Llama 2 7B. Parameters for the decoding strategies are tuned so that typical renormalizing quantitites $Z_{k},$ $Z_{\pi}$ and $Z_{\tau}$ are of the same size. We then plot the log of the ratio of the local normalization distortion for the pair of generated texts, and then plot the cumulative distribution function. We see in each case that nucleus sampling produces the smallest local normalization distortion, followed by temperature sampling and top-k sampling.
The experiments reported in the main body of the text were carried out using Llama 2 7B. We repeat these experiments here for other language models from the Llama and Pythia families, and find that results align. We tune values of p and $\tau$ to $k$ as described in Section 6.1. This process will have some noise and we note that the tuned values are slightly different for each model.
Figure 4: Figure 2 repeated for Llama 3.2 1B.
Figure 5: Figure 2 repeated for Llama 3.2 3B.
Figure 6: Figure 2 repeated for Pythia 1B.
Figure 7: Figure 2 repeated for Pythia 2.8B.
Figure 8: Figure 3 repeated for Llama 3.2 1B.
Figure 9: Figure 3 repeated for Llama 3.2 3B.
Figure 10: Figure 3 repeated for Pythia 1B.
Figure 11: Figure 3 repeated for Pythia 2.8B.
@@ Appendix F A Comment on Timescales for Normalization
$\frac{1}{C}$ $\leq\dfrac{q_{\tau}^{\prime\prime}(\{\underline{z}\in\mathcal{A}^{\infty}:z_{1}\cdots z_{T}=w_{1}\cdots w_{T}\})}{p^{\frac{1}{\tau}}(w_{1}\cdots w_{T})\exp(T.P)}$
$\leq C.$
@@
Appendix G A Note on Quality and Diversity
Quality and diversity are broad terms which could admit several interpretations. The notion that our metrics attempt to capture can be described as follows. Consider the scenario in which a language model is prompted ‘please tell me a joke’. In our interpretation, the model would be judged high on the quality metric even if it always responded with the same joke, provided the joke was a good one. The model would be judged high on the diversity metric provided there is a high chance that, when prompted twice, it would produce different outputs, even if these outputs are nonsense, or if an individual output repeats itself, or if the different outputs are semantically similar.
Appendix H Comparisons with the Article: Gareev et al. (2024)
We would like to thank one of our anonymous referees for making us aware of the article Gareev et al. (2024). Their headline conclusion, ‘in most configurations, global decoding performs worse than the local decoding versions of the same algorithms’, seems in direct opposition to ours, while their results are actually entirely consistent with our own. We carefully point out how the apparent inconsistencies arise.
H.1 Parameter Pinning and Results on Diversity
In comparing the effect of local and global decoding on (proxies for) quality and diversity, we follow the approach of Caccia et al. (2019) of plotting quality and diversity on the same plot (as in Figure 2). This allows us to compare, for example, locally normalized top-k vs globally normalized top-k across the whole range of parameters. Gareev et al. (2024) show that, for any fixed value of $k$, local decoding produces text which is more diverse than global decoding. We show that, at any desired level of quality of output, local decoding produces text which is less diverse than global decoding. There is no contradiction here and we agree with their result, as can be seen that the blue and red dots in Figure 2 are generally to the left of the orange and green dots. The apparent discrepancy arises since, if one wishes to compare local and global top-k at some fixed threshold for quality, one needs to use different values of k for the local and global decoding.
H.2 Results on Quality
As with diversity, Gareev et al. (2024) plot quality against parameter for local and global decoding, rather than plotting quality and diversity. However, in the case of their results on quality, we do not think this is the cause of the apparent discrepancy between our results and theirs. Instead, we think there are two fundamental factors.
MAUVE, their measure of quality, works by measuring both type 1 errors, where a language model produces un-human like text, and type-2 errors, where a language model fails to capture the full diversity of human language. This does not align with our desired notion of quality, which relates only to type-1 errors, see Section 4. Instead, in our language, MAUVE measures a convex combination of quality and diversity, and as such there is no clear discrepancy between their results and ours.
We suspect however that a much larger factor is at play. In our experiments we produce texts of constant length. Gareev et al. (2024) have a rather clever way of approximately sampling from the globally normalized distribution, which allows them to produce much longer texts. A consequence of not requiring fixed length generation is that their globally sampled texts are much shorter than their locally sampled ones, for some parameters by a factor of nearly 4. This length discrepancy is the first of their three suggested explanations of their results. The fact that global and local sampling produce outputs of such starkly differing lengths is very interesting, but isn’t really what we were hoping to measure when we talk about quality. Indeed, the fact that our generated texts are of constant length has allowed us to avoid the thorny question of whether we should be measuring quality, or quality per token, and we hope that further research might look at how our theoretical results would be affected if one were to normalize by generation length.
Appendix I Licenses
We have used Llama 2, Llama 3, and Pythia models under their respective community license agreements. [^2] See https://huggingface.co/meta-llama and https://huggingface.co/EleutherAI.. Use for research is consistent with the terms of these licenses.